please please answer this question...please
In the given equation
S2 O82-  + 2e-  ------> 2SO42-
Mn2+ +4H2O -------> MnO4+ 8H+ 5e-
how many moles of S2O82- ions are required to oxidize 1mole of Mn2+
and thanks in advance

Grade:12

1 Answers

Sahani Kumar
5 years ago
(Mn2+) + (H2O) ------->(MnO4-) + (8H+ )+ (5e-) 
so n=5
(S2O82-) + (2e-) --------->2SO4^2-      so n=2
These two reaction show that 
1 mole of Mn2+ yields ( 5e-)
but 1 mole S2O8^2- needs only 2e-
So now let's how many moles of S2O8^2- will accept 5e-    2e-  -------accepted by--------1 mole
                      5e- --------will be accepted by--------?mole
So ?mole =( 1mole * 5e- )/2e- =5/2 =2.5mole
 5e- will be accepted by 5/2 mole  =2.5 mole 
So answer is that 2.5 mole of S2O8^2- will accept 
5e- and therefore will oxidise 1 mole of Mn2+
Regards 
Askiitians members 
If you like this answer please encourage us by approving this answer.